OpenStudy (anonymous):

How do I solve this equation for x? (Will give medal & fan) 10 = 3^(x+5) - 2

OpenStudy (anonymous):

1) add 2 2) take natural lot 3) use log rule 4) divide ln(3) 5) subtract 5 :DDD

OpenStudy (anonymous):

And 4 isn't an answer option for this

OpenStudy (agent0smith):

@cutrightceo try plugging in x=4: 10 = 3^(4+5) - 2 10 = 3^9 - 2 10 = 19681 is 10 equal to 19,681?? Take logs of both sides (add 2 to both sides first): log12 = log3^(x+5) now you can bring down the exponent log12 = (x+5)log3

OpenStudy (anonymous):

so 1.079 = (x + 5)0.477

OpenStudy (anonymous):

I think I can solve it from here (: thank you!

OpenStudy (anonymous):

final answer I got is -2.738 (rounded to the nearest thousandth)

OpenStudy (anonymous):

Okay (: thank you